# A Computational Phase Function Approach for Obtaining $\alpha\alpha$ Wavefunctions

### v1

@@ 3 Results and Discussion
Figure 2: Wavefunction for s, p & d-waves at resonance energies in comparison with results of Hiura et. al.
@@
Figure 3: Wavefunction for s-wave at 19.45 MeV in comparison with results of Yoshiharu et. al.
@@ 4 Conclusion
The resulting wavefunctions exhibit physically consistent behavior over the entire radial domain, including smooth evolution in the interaction region and correct asymptotic oscillatory behavior at large distances. The angular-momentum dependence of the solutions is clearly manifested through the increasing influence of the centrifugal barrier from s- to g-waves, demonstrating that the method reliably captures the characteristic features of cluster–cluster scattering dynamics. The stabilization of the amplitude function at large radii further confirms the internal consistency of the approach and the suitability of the adopted interaction model.
A significant outcome of the present analysis is that the computed wavefunctions show very good agreement with previously reported results by Hiura et al., thereby providing an important validation of the present phase-function-based formulation. This agreement indicates that the essential physics of the $\alpha\alpha$ interaction is effectively encoded in the optimized potential and the phase evolution, even without explicitly invoking more elaborate microscopic formalisms. The consistency with established results strengthens confidence in the accuracy and reliability of the present computational approach.

### v2

@@ 3 Results and Discussion
Figure 2: (a) Scattering phase shifts and (b) interaction potentials for the $\ell=0$, 2, and 4 partial waves. The RPA results refer to the potentials constructed using the genetically optimized parameters reported by Sastri et al. [34].
Figure 3: Variation of the phase shift (in radians) as a function of distance (in fm) for the $\ell=0$, 2, and 4 partial waves at resonant energy. The RPA results correspond to the reference potential approach using the parameters reported by Sastri et al. [34].
Figure 4: Obtained scattering wavefunctions for S, P & D-waves in comparison with results of Hiura et. al. RPA refers to wavefunction obtained using reference potential approach parameters from Sastri et al. [34]
@@
Figure 5: Scattering wavefunction for S-wave at 19.45 MeV in comparison with results of Yoshiharu et al. [41]
@@ 4 Conclusion
The resulting wavefunctions exhibit physically consistent behavior over the entire radial domain, including smooth evolution in the interaction region and correct asymptotic oscillatory behavior at large distances. The angular-momentum dependence of the solutions is clearly manifested through the increasing influence of the centrifugal barrier from s- to g-waves, demonstrating that the method reliably captures the characteristic features of cluster-cluster scattering dynamics. The stabilization of the amplitude function at large radii further confirms the internal consistency of the approach and the suitability of the adopted interaction model.
A significant outcome of the present analysis is that the computed wavefunctions show very good agreement with previously reported results by Hiura et al., thereby providing an important validation of the present phase function based formulation. This agreement indicates that the essential physics of the $\alpha\alpha$ interaction is effectively encoded in the optimized potential and the phase evolution, even without explicitly invoking more elaborate microscopic formalisms. The consistency with established results strengthens confidence in the accuracy and reliability of the present computational approach.
@@
Appendix A Explicit Forms of the Amplitude and Wave Functions
For completeness, we provide below the fully expanded expressions used in the numerical implementation for the $\ell=0,2,$ and $4$ partial waves. The equations follow directly from the general phase–amplitude formalism presented in Sec. 2.
A.1 Amplitude Equations
The general amplitude equation is
$A_{\ell}^{\prime}(r)=-\frac{A_{\ell}(r)\,V(r)}{k\left(\hbar^{2}/2\mu\right)}\Big[\cos\delta_{\ell}(r)\,\hat{j}_{\ell}(kr)-\sin\delta_{\ell}(r)\,\hat{\eta}_{\ell}(kr)\Big]\Big[\sin\delta_{\ell}(r)\,\hat{j}_{\ell}(kr)+\cos\delta_{\ell}(r)\,\hat{\eta}_{\ell}(kr)\Big].$ (37)
Below we list the explicit forms.
$\ell=0$
Using
$\hat{j}_{0}(kr)=\sin(kr),\qquad\hat{\eta}_{0}(kr)=-\cos(kr),$
we obtain
$A_{0}^{\prime}(r)$ $=-\frac{A_{0}V(r)}{k\left(\hbar^{2}/2\mu\right)}\left[\cos\delta_{0}\sin(kr)+\sin\delta_{0}\cos(kr)\right]$
$\quad\times\left[\sin\delta_{0}\sin(kr)-\cos\delta_{0}\cos(kr)\right].$ (38)
$\ell=2$
For
$\hat{j}_{2}(kr)$ $=\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{kr}\cos(kr),$
$\hat{\eta}_{2}(kr)$ $=-\left(\frac{3}{(kr)^{2}}-1\right)\cos(kr)-\frac{3}{kr}\sin(kr),$
the amplitude equation becomes
$A_{2}^{\prime}(r)$ $=-\frac{A_{2}V(r)}{k\left(\hbar^{2}/2\mu\right)}\Big[\cos\delta_{2}\,\hat{j}_{2}(kr)-\sin\delta_{2}\,\hat{\eta}_{2}(kr)\Big]$
$\quad\times\Big[\sin\delta_{2}\,\hat{j}_{2}(kr)+\cos\delta_{2}\,\hat{\eta}_{2}(kr)\Big].$ (39)
$\ell=4$
For
$\hat{j}_{4}(kr)$ $=\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\left(\frac{10}{kr}-\frac{105}{(kr)^{3}}\right)\cos(kr),$
$\hat{\eta}_{4}(kr)$ $=-\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\cos(kr)+\left(\frac{10}{kr}-\frac{105}{(kr)^{3}}\right)\sin(kr),$
we obtain
$A_{4}^{\prime}(r)$ $=-\frac{A_{4}V(r)}{k\left(\hbar^{2}/2\mu\right)}\Big[\cos\delta_{4}\,\hat{j}_{4}(kr)-\sin\delta_{4}\,\hat{\eta}_{4}(kr)\Big]$
$\quad\times\Big[\sin\delta_{4}\,\hat{j}_{4}(kr)+\cos\delta_{4}\,\hat{\eta}_{4}(kr)\Big].$ (40)
A.2 Wave Function Expressions
The reconstructed reduced radial wave function is
$u_{\ell}(r)=A_{\ell}(r)\Big[\cos\delta_{\ell}(r)\,\hat{j}_{\ell}(kr)-\sin\delta_{\ell}(r)\,\hat{\eta}_{\ell}(kr)\Big].$ (41)
Explicitly:
$\ell=0$
$u_{0}(r)=A_{0}(r)\left[\cos\delta_{0}(r)\sin(kr)+\sin\delta_{0}(r)\cos(kr)\right].$ (42)
$\ell=2$
$u_{2}(r)=A_{2}(r)\Big[\cos\delta_{2}(r)\,\hat{j}_{2}(kr)-\sin\delta_{2}(r)\,\hat{\eta}_{2}(kr)\Big].$ (43)
$\ell=4$
$u_{4}(r)=A_{4}(r)\Big[\cos\delta_{4}(r)\,\hat{j}_{4}(kr)-\sin\delta_{4}(r)\,\hat{\eta}_{4}(kr)\Big].$ (44)